OpenStudy (lgbasallote):

Prove by contradiction that the sum of irrational and rational is irrational

OpenStudy (lgbasallote):

i suppose first step is to assume a and b to be the addends

OpenStudy (lgbasallote):

then i do a = 2m/4n and b = x/y

OpenStudy (lgbasallote):

then i assume the sum is rational

OpenStudy (lgbasallote):

\[\frac{2m}{4n} + \frac xy \implies \frac{2m + 4x}{4n + 4y}\]

OpenStudy (lgbasallote):

then... \[\frac{2(m + 2x)}{4(n+y)}\]

OpenStudy (lgbasallote):

this isn't simplified...so it's not rational

OpenStudy (lgbasallote):

is that how you do it?

OpenStudy (asnaseer):

let \(i_1\) be an irrational number and \(r_1, r_2\) be rational numbers. Then, if you assume the inverse statement, we get:\[i_1+r_1=r_2\]This leads to:\[i_1=r_2-r_1\]But \(r_2-r_1\) is a rational number. Therefore...

OpenStudy (lgbasallote):

i assume that means im wrong?

OpenStudy (asnaseer):

I couldn't follow your argument - where in your argument is the irrational number?

OpenStudy (lgbasallote):

2m/4n

OpenStudy (anonymous):

I'm confused about your assumptions. Why do you call a = \(\frac{2m}{4n}\)? Are m and n supposed to be integers? That would make a rational.

OpenStudy (asnaseer):

2m/4n does not represent an irrational number

OpenStudy (lgbasallote):

isn't rational number a number that's not in simplified form?

OpenStudy (anonymous):

No, a rational number cannot be expressed as a ratio of integers.

OpenStudy (asnaseer):

an irrational number is one that CANNOT be written as \(\displaystyle\frac{p}{q}\) where p and q are integers.

Parth (parthkohli):

@SmoothMath Perhaps you meant 'can'.

OpenStudy (lgbasallote):

but the proof of sqrt 2 is irrational involves using even/even

OpenStudy (anonymous):

Thanks, Parth. A rational number can. An irrational number cannot.

OpenStudy (asnaseer):

I think you may be confusing tis with the proof that \(\sqrt{2}\) is irrational where we start by assuming it IS rational and in its SIMPLIFIED form.

OpenStudy (lgbasallote):

yes i am confusing that

OpenStudy (asnaseer):

so 1/3 is a rational number. so is 12/24

OpenStudy (anonymous):

2/4 is rational.

Parth (parthkohli):

I see clarity in @asnaseer's proof. Cheers!

OpenStudy (asnaseer):

thanks @ParthKohli :)

OpenStudy (lgbasallote):

i see short steps though...

OpenStudy (asnaseer):

?

OpenStudy (anonymous):

Yeah, I agree with asnaseer. Now, for rigor, we could show that the difference of two rational numbers is rational. That's not hard to show though.

OpenStudy (lgbasallote):

how do you rewrite irrational number then?

OpenStudy (anonymous):

You cannot, lgba.

OpenStudy (asnaseer):

Yes @SmoothMath - that is why I ended with ... (to be completed by the asker of this question) :)

Parth (parthkohli):

The reason I find asnaseer's proof easier is that it doesn't involve the \(\Large {p\over q}\) stuff.

OpenStudy (anonymous):

To be a complete proof, it would have involved such kinds of things, although that part isn't too difficult.

OpenStudy (lgbasallote):

part of me really thinks this is too short....

OpenStudy (asnaseer):

@lgbasallote - you just need to represent an irrational number by /some/ symbol. I used \(i_1\) to represent it.

OpenStudy (asnaseer):

@lgbasallote - do you want me to complete the proof for you?

OpenStudy (lgbasallote):

rational - rational would be.. \[\frac ab - \frac xy\] \[\implies \frac {ay - bx}{by}\] so this is integer/integer

OpenStudy (lgbasallote):

i can manage @asnaseer

OpenStudy (anonymous):

Nailed it =)

OpenStudy (asnaseer):

yes - that is how to complete it. you have shown that \(r_2-r_1\) is rational.

Parth (parthkohli):

The contradiction is that we get \(\rm i_1 = r_3\) where \(\rm r_3 = r_2 - r_1\) and irrational cannot be rational.

OpenStudy (lgbasallote):

i envy the answerers of my questions though....they always get more than 3 medals....whilst i haven't gotten 3 medals in 1 question for a long time.....

OpenStudy (asnaseer):

@lgbasallote if you really want something to stick your teeth into, then try proving that:\[i_1^{i_2}\] can be rational. :) Here \(i_1, i_2\) are irrational numbers.

OpenStudy (lgbasallote):

i'm just in primary school...that is too advanced for me

Parth (parthkohli):

Thanks LG. You stole my identity. :P

OpenStudy (asnaseer):

@lgbasallote if you really are in primary school then I declare you a genius! :)

OpenStudy (lgbasallote):

why so?

OpenStudy (asnaseer):

Here in the UK, primary school means you are under 12 years old. And if you are doing problems like this at that age then you are indeed a genius!

OpenStudy (lgbasallote):

hmm...time to migrate to UK....

OpenStudy (asnaseer):

:)